OpenStudy (anonymous):

P(x)={-x-4, if x< -2 { 3/2x, if x≥ -2 Find domain and range and graph.

OpenStudy (anonymous):

Is it\[\frac{ 3 }{ 2 }x ..or,, \frac{ 3 }{ 2x }\] if it is \[\frac{ 3 }{ 2 }x\]then the domain is -infinity to plus infinity and the range -3 to infinity